invest a total of 16,000. part was savings paying 4% the rest was savings paying 6%. if the total interest for the year was $760, how much in each account?
Answer by stanbon(48557)

invest a total of 16,000. part was savings paying 4% the rest was savings paying 6%. if the total interest for the year was $760, how much in each account?
-------------
EQUATIONS:
Quantity: x + y = 16000
Interest: 0.04x + 0.06y = 760
-------------------
y = 16000-x
Substitute:
0.04x + 0.06(1-6000-x) = 760
4x + 6(16000-x) = 76000
-2x = 76000-6*16000
2x = 20000
x = $10,000 (amount invested at 4%)
-------------
16000-x = $6,000 (amount invested at 6%
------------
Cheers,
Stan H.